OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (johnweldon1993):

Okay same thing here....parallel lines have same slope....and use the point slope formula to solve for the equation y - (-4) = 3(x - 2) y + 4 = 3x - 6 y = 3x -10 and to check...plug in the 2 numbers -4 = 3(2) - 10 -4 = 6 - 10 6-6 yes so it works :)

OpenStudy (johnweldon1993):

Thanks for the medal :)